OpenStudy (anonymous):

In parallelogram LMNO, NO=10.2, and LO=14.7. What is the perimeter of parallelogram LMNO?

OpenStudy (anonymous):

HI Welcome to Openstudy! Perimeter is the sum of all sides of a polygon. So let's look at the parallelogram: Perimeter= LM+NO+LO+MN = 10.2 + 10.2 + 14.7 + 14.7 =?

OpenStudy (anonymous):

adding them up, what will you get?

OpenStudy (anonymous):

If we add that all up we get 49.8

OpenStudy (anonymous):

right (: so that will be your perimeter

OpenStudy (anonymous):

omg thank you that was help ful

OpenStudy (anonymous):

You're always welcome (: